OpenStudy (anonymous):

simplify 2x(3x+1)+2(x^2+3x+1)

OpenStudy (anonymous):

6x^2 +2x +2x^2 +6x+2=... can you finish it?

OpenStudy (anonymous):

\[2x(3x+1)+2(x^2+3x+1)\] \[=6x^2+2x+2x^2+6x+2\] \[=8x^2+8x+2\]